# Efficient Graph Optimization via Distance-Aware Graph Representation Learning

### v6

@@ 4.2 Main Theoretical Results
Our main theoretical contributions establish the following key results:
Theorem 1 (Generalization Bound of DRTR).
Assume the underlying ground-truth graph satisfies structural smoothness, and DRTR correctly removes $\epsilon$-fraction noisy edges and adds $\eta$-fraction semantically valid edges. Then, with high probability:
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+\mathcal{O}\left(\sqrt{\frac{|\mathcal{E}^{\prime}|\cdot\log|\mathcal{H}_{\text{DRTR}}|}{|\mathcal{V}_{L}|}}\right),$
where $|\mathcal{E}^{\prime}|\ll|\mathcal{E}|$ if DRTR successfully sparsifies the graph.
Theorem 2 (Convergence Rate).
Under standard assumptions (bounded gradients, Lipschitz continuity), the DRTR algorithm converges to a stationary point at rate $\mathcal{O}(1/\sqrt{T})$, where $T$ is the number of iterations.
Theorem 3 (Stability Guarantee).
@@
The detailed proofs of these theorems, along with comprehensive optimization analysis, computational complexity bounds, and additional theoretical properties, are provided in the Appendix.
@@ 4.3 Discussion
Our theoretical results support the intuition that DRTR improves both robustness and generalization by:
  • Reducing variance through selective pruning of noisy edges.
  • Enhancing representation power via topology completion.
  • Regularizing the learning objective through adaptive neighborhood smoothing.
The detailed mathematical analysis, including graph perturbation theory, noise reduction bounds, convergence guarantees, and comprehensive proofs, is provided in the Appendix.
@@ 5 Experimental Analysis
We conduct comprehensive experiments to evaluate the effectiveness of our proposed DRTR framework, including its modular variants: GDRA (Distance Recomputator), GKHDA (K-hop Diffusion Aggregator), and GKHDDRA (Full version). We address the following questions:
  • Q1:
    How does DRTR improve performance across diverse datasets and GNN backbones?
  • Q2:
    What is the contribution of each module (DR, TR, Diffusion)?
  • Q3:
    Can DRTR generalize to real-world downstream tasks such as recommendation or molecular modeling?
@@ 5.1 Datasets and Baselines
We evaluate on five benchmark datasets: Cora, Citeseer, Pubmed (standard citation graphs), ogbn-arxiv, and ogbn-products (from OGB Hu et al. (2020)). Each dataset varies in terms of scale, edge density, and semantic complexity.
@@
Model Cora Citeseer Pubmed ogbn-arxiv ogbn-products
Acc Time(s) Var Acc Time(s) Var Acc Time(s) Var Acc Acc
GCN 81.2 3.5 0.021 70.9 3.6 0.025 79.3 12.7 0.018 70.5 75.4
GCN+GDRA 82.6 3.9 0.016 71.3 3.8 0.019 80.1 12.5 0.016 73.1 76.6
GCN+GKHDA 82.4 4.0 0.017 71.7 4.1 0.018 80.5 12.4 0.016 72.8 76.1
GCN+GKHDDRA 82.7 4.2 0.013 72.3 4.3 0.014 80.9 12.3 0.014 73.9 77.2
SGC 74.2 2.1 0.030 71.5 2.3 0.026 78.2 10.4 0.024 68.2 74.1
SGC+GDRA 75.8 2.3 0.022 73.1 2.5 0.021 81.2 10.3 0.019 70.3 75.2
SGC+GKHDA 75.1 2.4 0.021 73.4 2.6 0.020 81.6 10.2 0.018 70.0 75.0
SGC+GKHDDRA 77.4 2.6 0.018 74.6 2.8 0.017 82.5 10.1 0.017 71.2 76.3
SSGC 83.0 2.5 0.015 75.6 2.6 0.017 73.6 11.3 0.016 69.9 74.8
SSGC+GDRA 83.2 2.6 0.013 76.4 2.7 0.015 74.2 11.1 0.014 71.1 75.4
SSGC+GKHDA 84.3 2.7 0.012 76.1 2.8 0.014 74.5 11.0 0.013 71.5 75.6
SSGC+GKHDDRA 84.1 2.8 0.011 77.6 2.9 0.013 74.7 10.9 0.012 72.0 76.8
APPNP 82.3 4.5 0.017 75.2 4.7 0.021 71.5 13.2 0.020 71.6 75.9
APPNP+GDRA 83.5 4.7 0.014 74.4 4.9 0.018 73.6 13.0 0.018 72.9 76.5
APPNP+GKHDA 83.8 4.9 0.013 74.5 5.0 0.016 74.1 12.9 0.016 73.1 76.3
APPNP+GKHDDRA 84.6 5.1 0.011 75.3 5.2 0.014 74.5 12.8 0.014 74.3 77.6
@@
Table 1 summarizes the results across five graph datasets and four representative GNN backbones. We observe several consistent trends:
  • Accuracy Gains: Across all datasets and models, DRTR variants outperform their corresponding baselines. The full version (GKHDDRA) achieves the best accuracy in nearly all settings. For example, SGC improves from 74.2% to 77.4% on Cora, and APPNP improves from 82.3% to 84.6%.
  • Stability Improvements: In all cases, DRTR-enhanced models exhibit lower performance variance across training runs, indicating more stable optimization dynamics and reduced sensitivity to initialization.
  • Low Overhead: While DRTR introduces additional computations (notably in the diffusion and reconstruction phases), the increase in training time remains modest. For instance, GCN’s training time on Pubmed grows from 12.7s to 12.3s (with DRTR), showing that dynamic reconstruction does not compromise scalability.
  • Generalization Across Models: Improvements hold across diverse architectures, including shallow models like SGC and diffusion-style models like APPNP and SSGC, suggesting that DRTR can be used as a generic plug-in for a wide range of GNNs.
@@
We perform ablation to understand the effect of each module:
  • GDRA (Distance Recomputator): removes noisy links and helps stabilize learning.
  • GKHDA (K-hop Heat Diffusion Aggregator): enhances long-range message passing.
  • GKHDDRA: combines the above with resampling and reconstruction for maximal performance.
Module Improvement $\uparrow$ Variance $\downarrow$
GDRA +1.4% 23.8%
GKHDA +1.2% 19.0%
TR (in full) +0.3% 18.8%
DRTR (full) +1.5% 38.1%
@@
To assess the broader applicability of DRTR, we extend our evaluation to two real-world downstream tasks beyond node classification:
@@
Model AUC AP
GraphSAGE 93.1 91.7
GraphSAGE+GDRA 94.0 92.5
GraphSAGE+GKHDA 94.3 92.8
GraphSAGE+GKHDDRA 95.1 93.6
@@
Model logP QED SA
GCN 0.423 0.218 0.387
GCN+GDRA 0.401 0.205 0.375
GCN+GKHDA 0.395 0.203 0.372
GCN+GKHDDRA 0.383 0.197 0.366
@@
These results show DRTR’s ability to:
  • Improve ranking and predictive performance in sparse user–item graphs (recommendation).
  • Preserve fine-grained local-global dependencies in dense, noisy molecular graphs.
This validates DRTR’s generality across structural domains—sparse bipartite graphs, dense chemical graphs, and traditional citation networks.
@@ 6 Conclusion
In this paper, we propose DRTR, a distance-aware graph representation learning framework that can optimize GNNs by dynamically refining multi-hop neighborhoods and topological structures. DRTR integrates a Distance Recomputator to prune noisy links based on semantic similarity, and a Topology Reconstructor to add meaningful long-range connections. Combined with a heat-diffusion-based message passing scheme, DRTR enables robust and expressive node representations. Extensive experiments across node classification, link prediction, and molecular property prediction confirm that DRTR consistently improves accuracy, stability, and generalization, while maintaining scalability across both synthetic and real-world graphs.
@@ Appendix A Appendix
A.1 Complete Proof of Theorem 1: Generalization Bound
Theorem 4 (Complete Generalization Bound of DRTR).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $|\mathcal{V}|=n$ nodes, and let $\mathcal{G}^{*}=(\mathcal{V},\mathcal{E}^{*})$ be the ground-truth graph satisfying structural smoothness. Assume:
  • The observed graph $\mathcal{G}$ contains $\epsilon$-fraction noisy edges
  • DRTR removes $\epsilon$-fraction noisy edges and adds $\eta$-fraction semantically valid edges
  • The GNN is $L$-Lipschitz with respect to input features
  • Node features are bounded: $\|\mathbf{x}_{v}\|_{2}\leq B$ for all $v\in\mathcal{V}$
@@ A.1 Complete Proof of Theorem 1: Generalization Bound
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+\mathcal{O}\left(\sqrt{\frac{|\mathcal{E}^{\prime}|\log n}{|\mathcal{V}_{L}|}}+\sqrt{\frac{\log(1/\delta)}{|\mathcal{V}_{L}|}}\right),$
where $|\mathcal{E}^{\prime}|=|\mathcal{E}|-\epsilon|\mathcal{E}|+\eta|\mathcal{E}|$ is the size of the reconstructed edge set.
Complete Proof.
We provide a detailed proof using Rademacher complexity analysis and structural graph properties.
@@
$\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})=\mathbb{E}_{\sigma}\left[\sup_{h\in\mathcal{H}_{\text{DRTR}}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}h(\mathbf{x}_{i})\right],$
@@
$|\mathcal{N}_{\text{eff}}(v)|=\sum_{k=1}^{K}|\mathcal{N}^{(k)}(v)\cap\{u:d_{vu}^{(k)}\leq\alpha_{k}\}|.$
The key insight is that $d_{\text{eff}}\ll d_{\text{original}}$ due to distance-based pruning.
@@
$\|f_{\text{DRTR}}(\mathbf{x}_{v})-f_{\text{DRTR}}(\mathbf{x}_{u})\|_{2}\leq L_{\text{DRTR}}\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2},$
@@
where $L_{\text{DRTR}}=L\cdot\prod_{k=1}^{K}(1+\lambda_{k})$ accounts for the multi-hop aggregation and distance recomputation.
Step 4: Complexity Bound
Using the covering number argument and the fact that DRTR reduces effective neighborhood size, we obtain:
@@
$\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})\leq\mathcal{O}\left(\sqrt{\frac{d_{\text{eff}}\log n}{n}}\right).$
Step 5: Final Bound
Combining the Rademacher complexity bound with standard generalization theory:
@@
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+2\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right).$

### v7

@@ 4.2 Main Theoretical Results
The theoretical results show that DRTR optimizes the computational graph by reducing noise propagation and improving connectivity, which are core primitives underlying many graph optimization problems. By tightening generalization bounds and improving stability under perturbations, DRTR provides theoretical guarantees for efficient graph optimization in large, noisy, or evolving graphs.
Our main theoretical contribution focuses on how DRTR’s adaptive neighborhood pruning reduces effective degree and improves generalization:
Theorem 4.1 (Generalization Bound under Adaptive Neighborhood Pruning).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $|\mathcal{V}|=n$ nodes, and let $d_{\text{eff}}$ denote the average effective degree after DRTR’s distance-based pruning. Assume the underlying ground-truth graph satisfies structural smoothness, and DRTR correctly removes $\epsilon$-fraction noisy edges and adds $\eta$-fraction semantically valid edges. Then, with probability at least $1-\delta$:
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+\mathcal{O}\left(\sqrt{\frac{d_{\text{eff}}\cdot\log n}{|\mathcal{V}_{L}|}}+\sqrt{\frac{\log(1/\delta)}{|\mathcal{V}_{L}|}}\right),$
where $d_{\text{eff}}=\frac{1}{n}\sum_{v\in\mathcal{V}}|\mathcal{N}_{\text{eff}}(v)|$ and $\mathcal{N}_{\text{eff}}(v)=\bigcup_{k=1}^{K}\{u\in\mathcal{N}^{(k)}(v):d_{vu}^{(k)}\leq\alpha_{k}\}$ is the effective neighborhood after pruning. The key insight is that $d_{\text{eff}}\ll d_{\text{original}}$ due to distance-based filtering, leading to reduced hypothesis complexity and improved generalization.
The detailed proof, which establishes the connection between effective degree reduction and Rademacher complexity bounds, is provided in the Appendix.
Proposition 4.2 (Convergence Rate).
Under standard assumptions (bounded gradients, Lipschitz continuity), and noting that DRTR’s distance recomputation and topology reconstruction modules preserve $L$-smoothness and bounded gradient assumptions, the DRTR algorithm converges to a stationary point at rate $\mathcal{O}(1/\sqrt{T})$, where $T$ is the number of iterations. The proof follows standard SGD analysis (see Appendix).
Proposition 4.3 (Stability under Graph Perturbation).
@@
where $C$ is a constant depending on the Lipschitz constant and graph properties. This stability guarantee explains DRTR’s robustness to noisy and dynamically evolving graphs (see Appendix for proof).
@@ 4.3 Discussion
DRTR improves robustness and generalization by reducing variance (selective pruning), enhancing representation power (topology completion), and regularizing the objective (adaptive neighborhood smoothing). Detailed proofs are in the Appendix.
@@ 5 Experimental Analysis
We evaluate DRTR from the perspective of graph optimization quality, where optimization refers to improving the structure, connectivity, and effective neighborhoods that downstream tasks rely on. We conduct comprehensive experiments to evaluate the effectiveness of our proposed DRTR framework. We evaluate three DRTR-based variants:
  • GDRA: baseline + Distance Recomputator (DR only, Section 3.2);
  • GKHDA: baseline + K-hop heat diffusion aggregator (Section 3.1) without DR/TR;
  • GKHDDRA: full DRTR with diffusion, DR, and TR (Sections 3.1–3.3).
We address three key questions: (Q1) How does DRTR improve performance across diverse datasets and GNN backbones? (Q2) What is the contribution of each module (DR, TR, Diffusion)? (Q3) Can DRTR generalize to real-world downstream tasks such as recommendation or molecular modeling?
@@ 5.1 Datasets and Baselines
We evaluate on five benchmark datasets: Cora, Citeseer, Pubmed (standard citation graphs), ogbn-arxiv, and ogbn-products (from OGB hu2020open). Each dataset varies in terms of scale, edge density, and semantic complexity.
@@
[…]
tab:all-results Model Cora Citeseer Pubmed ogbn-arxiv ogbn-products Acc Time(s) Var Acc Time(s) Var Acc Time(s) Var Acc Acc GCN 81.2 3.5 0.021 70.9 3.6 0.025 79.3 12.7 0.018 70.5 75.4 GCN+GDRA 82.6 3.9 0.016 71.3 3.8 0.019 80.1 12.5 0.016 73.1 76.6 GCN+GKHDA 82.4 4.0 0.017 71.7 4.1 0.018 80.5 12.4 0.016 72.8 76.1 GCN+GKHDDRA 82.7 4.2 0.013 72.3 4.3 0.014 80.9 12.3 0.014 73.9 77.2 SGC 74.2 2.1 0.030 71.5 2.3 0.026 78.2 10.4 0.024 68.2 74.1 SGC+GDRA 75.8 2.3 0.022 73.1 2.5 0.021 81.2 10.3 0.019 70.3 75.2 SGC+GKHDA 75.1 2.4 0.021 73.4 2.6 0.020 81.6 10.2 0.018 70.0 75.0 SGC+GKHDDRA 77.4 2.6 0.018 74.6 2.8 0.017 82.5 10.1 0.017 71.2 76.3 SSGC 83.0 2.5 0.015 75.6 2.6 0.017 73.6 11.3 0.016 69.9 74.8 SSGC+GDRA 83.2 2.6 0.013 76.4 2.7 0.015 74.2 11.1 0.014 71.1 75.4 SSGC+GKHDA 84.3 2.7 0.012 76.1 2.8 0.014 74.5 11.0 0.013 71.5 75.6 SSGC+GKHDDRA 84.1 2.8 0.011 77.6 2.9 0.013 74.7 10.9 0.012 72.0 76.8 APPNP 82.3 4.5 0.017 75.2 4.7 0.021 71.5 13.2 0.020 71.6 75.9 APPNP+GDRA 83.5 4.7 0.014 74.4 4.9 0.018 73.6 13.0 0.018 72.9 76.5 APPNP+GKHDA 83.8 4.9 0.013 74.5 5.0 0.016 74.1 12.9 0.016 73.1 76.3 APPNP+GKHDDRA 84.6 5.1 0.011 75.3 5.2 0.014 74.5 12.8 0.014 74.3 77.6
@@
[…]
fig:dataset-comparison Refer to caption
Figure 1: Performance across datasets. (a) Accuracy heatmap showing improvements across all model-dataset combinations. (b) Average improvement by dataset. (c) Variance reduction demonstrating improved stability. (d) Best performance comparison. (e) Training time overhead showing modest computational cost.
[…]
fig:comprehensive-performance provides a comprehensive performance analysis across multiple dimensions. Panel (a) shows accuracy improvements by model, where DRTR variants consistently outperform baselines across GCN, SGC, SSGC, and APPNP. Panel (b) demonstrates variance reduction, with GKHDDRA achieving up to 38.1% variance reduction on Cora. Panel (c) illustrates training time overhead, showing that DRTR introduces only modest computational overhead (4-20% depending on model). Panel (d) presents accuracy comparison across all five datasets using GCN backbone, demonstrating consistent improvements. Panel (e) shows component contributions, confirming that the full DRTR achieves the best performance.
We observe several consistent trends. Across all datasets and models, DRTR variants outperform their corresponding baselines, with the full version (GKHDDRA) achieving the best accuracy in nearly all settings. For example, SGC improves from 74.2% to 77.4% on Cora (+3.2%), and APPNP improves from 82.3% to 84.6% (+2.3%). On average, GKHDDRA achieves 1.5–3.4% accuracy improvement across datasets, with the largest gains on ogbn-arxiv (3.4%). In all cases, DRTR-enhanced models exhibit lower performance variance across training runs, indicating more stable optimization dynamics and reduced sensitivity to initialization. For instance, variance reduction ranges from 23.8% (GDRA on Cora) to 44% (GKHDDRA on Citeseer), with an average reduction of 38.1% for the full DRTR variant.
While DRTR introduces additional computations (notably in the diffusion and reconstruction phases), the increase in training time remains modest (4–20% depending on model). For instance, GCN’s training time on Pubmed decreases slightly from 12.7s to 12.3s (with DRTR), indicating that dynamic reconstruction does not introduce noticeable overhead and can even stabilize convergence (leading to fewer epochs to early stopping). Some models show improved training time due to faster convergence to optimal performance. Improvements hold across diverse architectures, including shallow models like SGC and diffusion-style models like APPNP and SSGC, suggesting that DRTR can be used as a generic plug-in for a wide range of GNNs.
@@
[…]
fig:comprehensive-performance Refer to caption
Figure 2: Comprehensive performance analysis. (a) Accuracy improvement by model showing consistent gains across GCN, SGC, SSGC, and APPNP. (b) Variance reduction demonstrating improved stability. (c) Training time overhead showing modest computational cost. (d) Accuracy across datasets using GCN backbone. (e) Component contribution analysis on Cora.
@@
[…]
Table LABEL:tab:component-mini quantifies the individual contributions: GDRA (Distance Recomputator) alone provides +1.4% accuracy improvement and 23.8% variance reduction on Cora, GKHDA (K-hop Heat Diffusion Aggregator) provides +1.2% improvement and 19.0% variance reduction, while TR contributes an additional +0.3% improvement and 18.8% variance reduction. The full DRTR achieves +1.5% improvement and 38.1% variance reduction, demonstrating that the modules work synergistically rather than additively. GDRA removes noisy links and helps stabilize learning, GKHDA enhances long-range message passing, and GKHDDRA combines the above with topology reconstruction for maximal performance.
[…]
fig:component-ablation Refer to caption
Figure 3: Component ablation analysis. (a) Accuracy improvement by component showing individual and combined contributions. (b) Variance reduction demonstrating stability improvements. (c) Combined effect analysis showing the optimal trade-off achieved by full DRTR.
[…]
tab:component-mini Module Improvement $\uparrow$ Variance $\downarrow$ GDRA +1.4% 23.8% GKHDA +1.2% 19.0% TR (in full) +0.3% 18.8% DRTR (full) +1.5% 38.1%
@@
To assess the broader applicability of DRTR, we extend our evaluation to two real-world downstream tasks beyond node classification. These two tasks—link prediction and molecular property prediction—are widely used as surrogate optimization benchmarks, because better graph representations lead to better optimization performance (ranking quality or regression accuracy):
@@
[…]
tab:recsys Model AUC AP GraphSAGE 93.1 91.7 GraphSAGE+GDRA 94.0 92.5 GraphSAGE+GKHDA 94.3 92.8 GraphSAGE+GKHDDRA 95.1 93.6
@@
[…]
tab:zinc Model logP QED SA GCN 0.423 0.218 0.387 GCN+GDRA 0.401 0.205 0.375 GCN+GKHDA 0.395 0.203 0.372 GCN+GKHDDRA 0.383 0.197 0.366
@@
These results demonstrate DRTR’s ability to improve ranking and predictive performance in sparse user–item graphs (recommendation) while preserving fine-grained local-global dependencies in dense, noisy molecular graphs. On MovieLens-100K, DRTR achieves relative improvements of +2.0% AUC (from 93.1% to 95.1%) and +1.9% AP (from 91.7% to 93.6%) over the baseline GraphSAGE. On ZINC-12K, DRTR reduces MAE by 9.5% for logP (from 0.423 to 0.383), 9.6% for QED (from 0.218 to 0.197), and 5.4% for SA (from 0.387 to 0.366), demonstrating consistent improvements across different molecular properties.
This validates DRTR’s generality across structural domains—sparse bipartite graphs, dense chemical graphs, and traditional citation networks—with relative performance gains ranging from 2.0% to 9.6% depending on the task and metric.
@@ 6 Conclusion
In this paper, we propose DRTR, a distance-aware graph representation optimization framework that can optimize GNNs by dynamically refining multi-hop neighborhoods and topological structures. DRTR integrates a Distance Recomputator to prune noisy links based on semantic similarity, and a Topology Reconstructor to add meaningful long-range connections. Combined with a heat-diffusion-based message passing scheme, DRTR enables robust and expressive node representations. Extensive experiments across node classification, link prediction, and molecular property prediction confirm that DRTR consistently improves accuracy, stability, and generalization, while maintaining scalability across both synthetic and real-world graphs. Taken together, DRTR provides an efficient and general optimization layer that improves graph structure, optimizes message-passing pathways, and enhances the solution quality of downstream graph optimization problems.
@@ Appendix A Appendix
A.1 Complete Proof of Theorem 1 (Main Text): Generalization Bound
Theorem A.1 (Complete Generalization Bound under Adaptive Neighborhood Pruning).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $|\mathcal{V}|=n$ nodes, and let $d_{\text{eff}}$ denote the average effective degree after DRTR’s distance-based pruning. Let $\mathcal{G}^{*}=(\mathcal{V},\mathcal{E}^{*})$ be the ground-truth graph satisfying structural smoothness. Assume the observed graph $\mathcal{G}$ contains $\epsilon$-fraction noisy edges, DRTR correctly removes $\epsilon$-fraction noisy edges and adds $\eta$-fraction semantically valid edges, the GNN is $L$-Lipschitz with respect to input features, and node features are bounded such that $\|\mathbf{x}_{v}\|_{2}\leq B$ for all $v\in\mathcal{V}$.
@@ A.1 Complete Proof of Theorem 1 (Main Text): Generalization Bound
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+\mathcal{O}\left(\sqrt{\frac{d_{\text{eff}}\cdot\log n}{|\mathcal{V}_{L}|}}+\sqrt{\frac{\log(1/\delta)}{|\mathcal{V}_{L}|}}\right),$
where $d_{\text{eff}}=\frac{1}{n}\sum_{v\in\mathcal{V}}|\mathcal{N}_{\text{eff}}(v)|$ and $\mathcal{N}_{\text{eff}}(v)=\bigcup_{k=1}^{K}\{u\in\mathcal{N}^{(k)}(v):d_{vu}^{(k)}\leq\alpha_{k}\}$ is the effective neighborhood after pruning.
Proof A.2 (Complete Proof).
We provide a detailed proof using Rademacher complexity analysis and structural graph properties. The proof proceeds in several steps, establishing key lemmas that control error propagation through the multi-hop aggregation structure.
Notation. We denote by $I[A]$ the indicator function for event $A$. For each node $v\in\mathcal{V}$ and hop $k\in[K]$, let $t=N_{k}(v)$ denote the number of times node $v$ has been processed at hop $k$ up to the current iteration. We denote by $\mathbf{z}_{v}^{t}$ the representation of node $v$ after $t$ updates at hop $k$.
@@
$\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})=\mathbb{E}_{\sigma}\left[\sup_{h\in\mathcal{H}_{\text{DRTR}}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}h(\mathbf{x}_{i})\right],$ (17)
@@
$|\mathcal{N}_{\text{eff}}(v)|=\sum_{k=1}^{K}|\mathcal{N}^{(k)}(v)\cap\{u:d_{vu}^{(k)}\leq\alpha_{k}\}|.$ (18)
The key insight is that $d_{\text{eff}}\ll d_{\text{original}}$ due to distance-based pruning. Specifically, by the percentile-based thresholding mechanism, we have:
$d_{\text{eff}}=\frac{1}{n}\sum_{v\in\mathcal{V}}|\mathcal{N}_{\text{eff}}(v)|\leq(1-p)\cdot d_{\text{original}},$ (19)
where $p$ is the pruning percentile parameter.
@@
$\|f_{\text{DRTR}}(\mathbf{x}_{v})-f_{\text{DRTR}}(\mathbf{x}_{u})\|_{2}\leq L_{\text{DRTR}}\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2},$ (20)
@@
where $L_{\text{DRTR}}=L\cdot\prod_{k=1}^{K}(1+\lambda_{k})$ accounts for the multi-hop aggregation and distance recomputation. The Lipschitz constant can be bounded as:
@@
$L_{\text{DRTR}}\leq L\cdot\left(1+\frac{\lambda_{0}}{H}\right)^{K}\leq L\cdot e^{\lambda_{0}K/H}=\mathcal{O}(L),$ (21)
since $\lambda_{k}=\lambda_{0}\exp(-\rho k)+\lambda_{\min}$ decays exponentially with $k$.
Step 4: Covering Number and Complexity Bound
Using the covering number argument and the fact that DRTR reduces effective neighborhood size from $d_{\text{original}}$ to $d_{\text{eff}}$, we obtain:
$\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})\leq\mathcal{O}\left(\sqrt{\frac{d_{\text{eff}}\log n}{n}}\right),$ (22)
where the covering number $\mathcal{N}(\mathcal{H}_{\text{DRTR}},\epsilon)$ satisfies:
@@
$\log\mathcal{N}(\mathcal{H}_{\text{DRTR}},\epsilon)\leq d_{\text{eff}}\cdot\log\left(\frac{L_{\text{DRTR}}B}{\epsilon}\right),$ (23)
and $B$ is the bound on node features.
Step 5: Final Bound via Standard Generalization Theory
Combining the Rademacher complexity bound with standard generalization theory (see, e.g., bartlett2002rademacher):
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+2\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right).$ (24)